OpenStudy (andijo76):

i figured out the first part of my assignment but I am not sure if it is correct i will post the question and my answer if someone will see if I am on the right track

OpenStudy (andijo76):

Part I: Determine whether an individual's political affiliation affects his or her overall satisfaction with the economy. Analyze any potential differences by conducting the appropriate type of t-test. Be sure to use the proper df and show all the steps of calculations, including the test statistic. Include an explanation of how the critical value was determined. Explain why you chose that particular t-test.

OpenStudy (andijo76):

this is the information from my mini tab Test of mu = 70 vs not = 70 Variable N Mean StDev SE Mean 95% CI T satisfaction with the ec 6 3.500 1.378 0.563 (2.053, 4.947) -118.17 Variable P satisfaction with the ec 0.000

OpenStudy (andijo76):

so would i write it like this hypotheses Ho: mu=70 [H1:\mu \neq70\] df=n-1 =6-1=5 df=5 t= this is what I got for part 1